the sum of two numbers is 62. their difference is 16. find the numbers.

Answer by Maths68(1474)

Let x and y are two numbers
x+y=62 .....(1)
x-y=16......(2)

Add (1) and (2)
x+y=62 .....(1)
x-y=16......(2)
--------
2x=78
x=78/2
x=39
Put the value of x in (1) or (2) we have
39+y=62
y=62-39
y=23
Numbers are 39 and 23
Check
x-y=16
39-23=16
16=16
and
x+y=62
39+23=62
62=62